OpenStudy (anonymous):

Given that f(x) = x2 − 4x − 3 and g(x) = the quantity of x plus three, over four , solve for f(g(x)) when x = 9. @Teddyiswatshecallsme

OpenStudy (teddyiswatshecallsme):

:/ Sorry i can't help i havee no idea in functions....

OpenStudy (anonymous):

Below are two different functions, f(x) and g(x). What can be determined about their slopes? f(x) Alex earns 1600 dollars in 400 hours. x g(x) 1 −8 5 12 9 32 HOW ABOUT THIS ONE? @Teddyiswatshecallsme

OpenStudy (teddyiswatshecallsme):

Nope lol sorry....

OpenStudy (anonymous):

Do you know anyone that can help me ? @Teddyiswatshecallsme

OpenStudy (tkhunny):

\(Given\;that\;f(x) = x^{2} − 4x − 3\;and\;g(x) = \dfrac{x+3}{4},\;find\;f(g(x))\;for\;x=9.\) You should not be struggling with this. We're just evaluating functions. g(9) = (9+3)/4 = 12/4 = 3 f(3) = 3^2 - 4(3) - 3 = 9 - 12 - 3 = -3-3 = -6 Is there a part of that you actually can't do?

OpenStudy (anonymous):

I did g(9)=9+3/3 g(9)=3 and then I plug in f(3)x^2-4x-3 equaling out to 0... idk what I did wrong.

OpenStudy (anonymous):

below are two different functions, f(x) and g(x). What can be determined about their y-intercepts? f(x) = x + 4 (x) g(x) −1 8 1 0 2 −4 I also have a lot of trouble finding the y intercept when they give me a x and the g(x)

OpenStudy (anonymous):

 can you please help me @tkhunny